The manager of a sporting goods store uses cost-plus pricing to determine the profit-maximizing price of bicycles. The cost of a bicycle to the store is $80. The manager estimates that the price elasticity of demand is –3.0. What is the profit-maximizing price?
1
Expert's answer
2020-11-25T11:52:13-0500

firm is profit maximizing, then we know that MR=MC

"MR=P(1+\frac{1}{\epsilon})"

МС=МК=80

"80=P(1+\frac{1}{-3})"

P=120